### v3

# A new characterization of (pre)liminary C*-algebras

Abstract.
Given an arbitrary countable ordinal $\alpha$, we introduce the notion of type $I_{\alpha}$ C*-algebra and $\alpha$-subhomogeneous C*-algebra. When $\alpha=0$, these recover the notions of Fell C*-algebra and of commutative C*-algebra, respectively. When $\alpha=n<\omega$, these recover the notions of type $I_{n}$ C*-algebra and of $n$-subhomogeneous C*-algebra, respectively. We prove that a separable C*-algebra is liminary if and only if it is type $I_{\alpha}$ for some $\alpha<\omega_{1}$, and it is preliminary (i.e., has no infinite-dimensional irreducible representation) if and only if it is $\alpha$-subhomogeneous for some $\alpha<\omega_{1}$. We also prove that for any countable ordinal $\alpha$ there exists a separable C*-algebra that is type $I_{\alpha}$ and not type $I_{\beta}$ for $\beta<\alpha$, and a separable C*-algebra that is $\alpha$-subhomogeneous and not $\beta$-subhomogeneous for any $\beta<\alpha$.
Key words and phrases:
C*-algebra, type I C*-algebra, Fell algebra, Fell’s condition, liminary C*-algebra, postliminary C*-algebra, preliminary C*-algebra, type $I_{\alpha}$ C*-algebra, $\alpha$-subhomogeneous C*-algebra, descriptive set theory, Fell compactification.
2000 Mathematics Subject Classification:
Primary 46L05, 46L35; Secondary 03E15, 54D35
The author was partially supported by the Marsden Fund Fast-Start Grant VUW1816 and the Rutherford Discovery Fellowship VUW2002 “Computing the Shape of Chaos” from the Royal Society of New Zealand, the Starting Grant 101077154 “Definable Algebraic Topology” from the European Research Council, the Gruppo Nazionale per le Strutture Algebriche, Geometriche e le loro Applicazioni (GNSAGA) of the Istituto Nazionale di Alta Matematica (INDAM), and the University of Bologna. Part of this work was done during a visit of the author to Chalmers University of Technology and the University of Göthenburg. The hospitality of these institutions is gratefully acknowledged.
1. Introduction
This paper is a contribution to the study of the structure and classification of separable C*-algebra. We focus on the class of so-called liminary C*-algebras, also known as liminal or CCR. These are precisely the C*-algebras whose irreducible representations on a Hilbert space have the property that their image coincides with the algebra of compact operators. The more generous requirement that the image contains the algebra of compact operators yields the notion of postliminary C*-algebra, also known as postliminal or Type I. Within the class of postliminary C*-algebras, liminary C*-algebras are precisely those whose primitive spectrum endowed with the Jacobson topology has closed points—i.e., every prime ideal is maximal among the class of prime ideals, and in fact within the class of all ideals [27, Corollary 2.6].
The study of liminary and postliminary C*-algebras goes back to the early days of C*-algebra theory. Its forefathers are Kaplansky [33, 32] and Mackey [42, 39, 40, 41], who pioneered the study of liminary and postliminary C*-algebras in the 1950s. Their motivation included the classification problem for C*-algebras and their irreducible representations. The classification problem of irreducible representations of locally compact Hausdorff topological groups can be seen as a particular instance of the one of C*-algebras, by considering the corresponding group C*-algebra. The theory of liminary and postliminary C*-algebras was further developed in the 1960s by Dixmier [13, 14, 12, 16], Effros [19, 20], Fell [23, 25, 24], and Glimm [28] among others.
An upshot of these works is Glimm’s Theorem from [28] characterizing separable postliminary C*-algebras in terms of the corresponding classification problem for irreducible representations: a separable C*-algebra is postliminary if and only if its irreducible representation are concretely classifiable up to unitary equivalence, and the corresponding Borel structure induced on the spectrum of the C*-algebra is standard.
Since then, the study of the representation theory of C*-algebra has divided into two streams, divided by the type $I$ versus non-type $I$ dichotomy. In the non-type $I$ case, the study of the classification problem for irreducible C*-algebras and their “definable spectra” has refined the analysis from the perspective of Borel complexity theory [35, 52, 22, 31]. In the type $I$ case, more stringent notions have been introduced in the attempt to obtain a complete description of the C*-algebras under considerations and their spectra.
Among these, the notion of type $I_{0}$ C*-algebra, also known as Fell algebra, has been one of the most fruitful. Defined in terms of the so-call Fell condition, Fell algebras have been studied by several authors [49, 4, 1, 51]. Fell’s condition plays a crucial role in the characterization and classification of continuous trace C*-algebras, achieved by Dixmier and Douady in terms of bundles of elementary C*-algebras [15, 18]; see also [50]. More recently, the Dixmier–Douady classification has been extended by means of suitable groupoid models to arbitrary Fell algebras by an Huef, Kumjian, and Sims [2]; see also [9, 11].
Fell algebras admit several equivalent characterizations. Very recently, Enders and Shulman proved Fell’s condition to be equivalent to commutativity of the central sequence algebra [21]. Motivated by this characterization, they introduced Fell’s condition of order $k$ for a given positive integer $k$ (the original Fell condition corresponding to the case $k=0$), and proved it to be equivalent to $k$-subhomogeneity of the central sequence algebra [21].
In this work, we introduce a further generalization, by replacing $k$ with an arbitrary countable ordinal $\alpha$. We define the (ordinal-valued) Fell rank of an irreducible representation of a C*-algebra, capturing Fell’s conditions of higher order. The class of type $I_{\alpha}$ C*-algebras is defined by requiring that all their irreducible representations have Fell rank at most $1+\alpha$. For $\alpha=0$ this subsumes the classical notion of type $I_{0}$ C*-algebra, and for $\alpha<\omega$ one recovers the Enders–Shulman definition.
As observed in [21], a type $I_{0}$ and, more generally, a type $I_{n}$ C*-algebra is necessarily liminary, and the same holds for arbitrary countable ordinals. It is also remarked in [21] that there exists liminary C*-algebras that are not type $I_{n}$ for any $n<\omega$. We prove that by considering arbitrary countable ordinal one obtains a hierarchy that is complete, and includes all liminary C*-algebras.
Theorem 1.1.
Let $A$ be a separable C*-algebra. Then $A$ is liminary if and only if it is type $I_{\alpha}$ for some countable ordinal $\alpha$.
We also prove that the hierarchy is strict, namely, for any countable ordinal $\alpha$ there exist liminary C*-algebras that are type $I_{\alpha}$ and not type $I_{\beta}$ for any $\beta<\alpha$. Such C*-algebras can be chosen to be scattered, namely have countable spectrum, and no infinite-dimensional irreducible representations.
Fell algebras admit yet another characterization, isolated by Pedersen, in terms of abelian elements. An element of a C*-algebra is abelian if the hereditary C*-algebra it generates is abelian. Enders and Shulman extended this characterization to arbitrary type $I_{n}$ C*-algebras, by replacing abelian elements with elements with global rank at most $n$. This notion is obtained by demanding that the hereditary C*-algebra they generate be $n$-subhomogeneous. (A C*-algebra is $n$-subhomogeneous if its irreducible representations have dimension at most $n$. For $n=1$, this recovers the notion of abelian C*-algebra.)
We extend the Enders–Shulman characterization to type $I_{\alpha}$ C*-algebras for an arbitrary countable ordinal $\alpha$. This is obtained in terms of the (ordinal-valued) Pedersen rank of an element of a C*-algebra, generalizing the notion of abelian element (corresponding to the rank 0 case) and global rank $n$ (corresponding to the case of finite ordinals). In turn, this yields a corresponding notion of $\alpha$-subhomogeneous C*-algebra for an arbitrary countable ordinal $\alpha$. Whereas $n$-subhomogeneous C*-algebras have all irreducible representations of dimension at most $n$, when $\alpha$ is infinite an $\alpha$-subhomogeneous C*-algebra can have irreducible representations of arbitrary large, albeit finite, dimension. They are defined by requiring that all their elements have Pedersen rank at most $\alpha$.
We define a C*-algebra to be preliminary if it has no infinite-dimensional representations. A preliminary C*-algebra is, in particular, liminary. This class of C*-algebras has received attention recently, in the work of Courtney and Shulman [10]; see also [30]. As in the case of type $I_{\alpha}$ C*-algebras within liminary C*-algebras, we prove that the hierarchy of $\alpha$-subhomogeneous C*-algebras within preliminary C*-algebras is complete.
Theorem 1.2.
Let $A$ be a separable C*-algebra. Then $A$ is preliminary, i.e., has no infinite-dimensional irreducible representation, if and only if it is $\alpha$-subhomogeneous for some countable ordinal $\alpha$.
Again, this hierarchy is strict, as for any countable ordinal $\alpha$ one can find $\alpha$-subhomogeneous C*-algebras that are not $\beta$-subhomogeneous for any $\beta<\alpha$. Furthermore, one can even find such examples to be scattered. In fact, examples of such C*-algebras were previously considered by Lazar and Taylor [37].
Acknowledgments
We are thankful to Jeffrey Bergfalk, Luigi Caputi, Alessandro Codenotti, Eusebio Gardella, Ilja Gogic, Ivan Di Liberti, Aristotelis Panagiotopoulos, Ebrahim Samei, Tatiana Shulman, and Joseph Zielinski for several helpful conversations.
2. Some C*-algebra notions
In this section we recall some notions from C*-algebra theory, and particularly the notions of liminary and postliminary C*-algebra; see also [48, 17, 7].
2.1. Spectrum of a C*-algebra
[…]
$\mathrm{hull}\left(I\right):=\left\{t\in\mathrm{Prim}\left(A\right):I\subseteq t\right\}$
where $I$ varies among the closed ideals of $A$. (It follows from the fact that every closed ideal of $A$ is intersection of primitive ideals that such a topology is well-defined.) Such a topology is compact when $A$ is unital.
There is a canonical map $\hat{A}\rightarrow\mathrm{Prim}\left(A\right)$ sending a unitary equivalence class of irreducible representations of $A$ to the kernel of any of its representatives. One considers $\hat{A}$ as a topological space with respect to the topology that makes such a map open and continuous.
If $I$ is a closed ideal of $A$, then the map
$t\mapsto t\cap I$
establishes a homeomorphism
$\mathrm{Prim}\left(A\right)\setminus\mathrm{hull}\left(I\right)\rightarrow\mathrm{Prim}\left(I\right)$
while the map
$t\mapsto t/I$
establishes a homeomorphism
$\mathrm{hull}\left(I\right)\rightarrow\mathrm{Prim}\left(A/I\right)\text{;}$
see [48, Theorem 4.1.11]. The assignment
$I\mapsto\mathrm{Prim}\left(A\right)\setminus\mathrm{hull}\left(I\right)\cong\mathrm{Prim}\left(I\right)$
establishes an order isomorphism between the lattice of closed ideals in $A$ and the lattice of open subsets of $\mathrm{Prim}\left(A\right)$.
Let $A$ be a separable C*-algebra. The quasi-state space $Q\left(A\right)$ of $A$ is the space of positive linear functionals on $A$ of norm at most $1$, which is a w*-closed subspace of the dual of $A$ [48, Section 3.2.1]. The space $P\left(A\right)$ of non-zero extreme points of $Q\left(A\right)$ is a $G_{\delta}$ subspace of $Q\left(A\right)$, whence Polish when endowed with the subspace topology [48, Proposition 4.3.2]. The elements of $P\left(A\right)$ are called pure states of $A$.
Let $A$ be a separable C*-algebra. The Borel T-structure (T stands for topological) on $\hat{A}$ is the $\sigma$-algebra generated by the open sets in the Jacobson topology, while the Borel M-structure (M stands for Mackey) on $\hat{A}$ is the one induced from the standard Borel structure on $P\left(A\right)$ by the map $P\left(A\right)\rightarrow\hat{A}$ mapping a pure state $\phi$ to the kernel of the corresponding irreducible representation $\pi_{\phi}$ [48, Theorem 4.3.3]. Then we have that the Borel T-structure is weaker than the Borel M-structure, but they coincide (and they are standard) when $A$ is type $I$ [48, Proposition 6.3.2].
Suppose that $x\in A$. Then $x$ defines a lower semi-continuous function $\check{x}:\mathrm{Prim}\left(A\right)\rightarrow\mathbb{R}$ by $\check{x}\left(J\right):=\left\|x+J\right\|_{A/J}$. In fact, the Jacobson topology on $\mathrm{Prim}\left(A\right)$ is the weakest topology that makes $\check{x}$ lower semi-continuous for all $x$ in (some countable dense subset of) $A$; see [48, Section 4.4].
If $A$ and $B$ are separable C*-algebras, and at least one between $A$ and $B$ is type $I$, then there exists a canonical continuous map from $\hat{A}\times\hat{B}$ to the spectrum of $A\otimes B$ [7, IV.3.4.22]. This is in fact a homeomorphism [7, IV.3.4.28], and it induces a homeomorphism
$\mathrm{Prim}\left(A\right)\times\mathrm{Prim}\left(B\right)\rightarrow\mathrm{Prim}\left(A\otimes B\right)\text{.}$
In particular, $\mathrm{Prim}\left(A\right)$ and $\mathrm{Prim}\left(A\otimes K\left(H\right)\right)$ are homeomorphic, since $\mathrm{Prim}\left(K\left(H\right)\right)$ is a singleton [7, IV.1.2.2].
2.2. Liminary and postliminary C*-algebras
Let $A$ be a separable C*-algebra, and $\pi$ be an irreducible representation of $A$ on a Hilbert space $H$. Then $\pi$ is CCR if $\pi\left(A\right)\subseteq K\left(H\right)$ (which implies $\pi\left(A\right)=K\left(H\right)$), and GCR if $\pi\left(A\right)\cap K\left(H\right)\neq 0$ (and hence $K\left(H\right)\subseteq\pi\left(A\right)$); see [7, IV.1.3.1]. A C*-algebra is CCR (respectively, GCR) if every irreducible representation of $A$ is CCR (respectively, GCR). A C*-algebra is liminary if and only if it is CCR; see [7, IV.1.3.1]. Recall also that an element $x$ of $A$ is abelian if the hereditary subalgebra $\overline{xAx^{\ast}}$ it generates is commutative [7, IV.1.1.1]. A Fell C*-algebra $A$ (also called type $I_{0}$ C*-algebra) is the C*-subalgebra generated by the abelian elements is equal to $A$ [7, IV.1.1.6]. Every Fell C*-algebra is liminary [7, IV.1.3.2]. A separable C*-algebra is elementary if it is isomorphic to $K\left(H\right)$ for some separable Hilbert space $H$ [7, IV.1.2.1]. Liminary C*-algebras were originally introduced by Kaplansky in [32]; see [48, Section 6.2.13]. Fell C*-algebras were introduced by Pedersen [48, Section 6.1.14], and can be characterized in terms of the so-called Fell condition [7, IV.1.4.17].
Definition 2.1.
Let $A$ be a separable C*-algebra and $\mathcal{C}$ be a class of C*-algebras. A decomposition series with subquotients in $\mathcal{C}$ for $A$ is a chain $\left(R_{\alpha}\right)_{\alpha<\omega_{1}}$ of closed ideals of $A$ such that:
  1. (1)
    $R_{0}=\left\{0\right\}$;
  2. (2)
    $R_{\alpha+1}/R_{\alpha}\in\mathcal{C}$ for every $\alpha<\omega_{1}$;
  3. (3)
    the union of $R_{\alpha}$ for $\alpha<\lambda$ is dense in $R_{\lambda}$ for every limit ordinal $\lambda$;
  4. (4)
    $R_{\alpha}=A$ eventually.
If $A$ is a separable C*-algebra, then it has a largest liminary closed ideal (i.e., a largest element in the collection of closed ideals that are liminal as C*-algebras); see [7, IV.1.3.9]. This can be explicitly defined as the ideal $J_{1}\left(A\right)$ comprising the $x\in A$ such that for every irreducible representation $\pi$ of $A$, $\pi\left(x\right)$ is a compact operator. Then one can recursively define a chain $\left(J_{\alpha}\left(A\right)\right)_{\alpha\leq\omega_{1}}$ of closed ideals, continuous at limits, such that
$J_{\alpha+1}\left(A\right)/J_{\alpha}\left(A\right)=J_{1}\left(A/J_{\alpha}\left(A\right)\right)\text{.}$
Let us also set $J_{0}\left(A\right)=0$. It is a feature of this definition that
$J_{\beta}\left(A\right)/J_{\alpha}\left(A\right)=J_{\beta}\left(A/J_{\alpha}\left(A\right)\right)$
for $\alpha\leq\beta$. The C*-algebra $A$ is postliminary if $A=J_{\alpha}\left(A\right)$ for some $\alpha<\omega_{1}$; see [7, IV.1.3.9].
A separable C*-algebra also has a largest Fell ideal [7, IV.1.1.8]. This can be explicitly defined as the C*-subalgebra $I_{1}\left(A\right)$ of $A$ generated by its abelian elements. (In fact, it is also the closed linear span of the abelian elements.) Again, one can recursively define the ideals $I_{\alpha}\left(A\right)$ for $\alpha\leq\omega_{1}$, and again $A$ is type $I$ if and only if $A=I_{\alpha}\left(A\right)$ for some $\alpha<\omega_{1}$ [7, IV.1.1.12]. Furthermore, $I_{\omega_{1}}\left(A\right)=J_{\omega_{1}}\left(A\right)$ is the largest type $I$ C*-subalgebra of $A$ [7, IV.1.1.12 and IV.1.3.9].
Let $A$ be a separable C*-algebra. A positive element $x$ of $A$ defines a lower semi-continuous function
$\hat{x}:\mathrm{Prim}\left(A\right)\rightarrow\left[0,\infty\right]\text{, }[\mathrm{ker}\left(\pi\right)]\mapsto\mathrm{Tr}\left(\pi\left(x\right)\right)\text{;}$
see [7, IV.1.4.8]. The element $x$ has continuous trace if $\hat{x}$ takes values in $[0,n]$ for some $n\in\mathbb{N}$ and it is continuous. Then the set $\mathfrak{m}\left(A\right)_{+}$ of positive continuous trace elements of $A$ is the positive part of an ideal $\mathfrak{m}\left(A\right)$ [7, IV.1.4.11 ]. A C*-algebra $A$ has continuous trace if $\mathfrak{m}\left(A\right)$ is dense in $A$.
Proposition 2.2.
Let $A$ be a separable C*-algebra. The following assertions are equivalent:
  1. (1)
    AA is postliminary;
  2. (2)
    AA is GCR;
  3. (3)
    AA is type $I$;
  4. (4)
    the canonical map $\hat{A}\rightarrow\mathrm{Prim}\left(A\right)$ is a homeomorphism;
  5. (5)
    the Borel T-structure on $\hat{A}$ is standard;
  6. (6)
    the Borel M-structure on $\hat{A}$ is standard;
  7. (7)
    AA admits a decomposition series with continuous trace subquotients.
Furthermore, the following conditions are equivalent:
  1. (a)
    AA is type $I$ and $\hat{A}$ is $T_{1}$;
  2. (b)
    AA is type $I$ and $\mathrm{Prim}\left(A\right)$ is $T_{1}$;
  3. (c)
    AA is liminary.
Proof.
The equivalence of (1)—(7) is the content of [7, IV.1.5.7, IV.1.5.12] and [48, Theorem 6.9.7]; see also [7, IV.1.5.7]. The equivalence of (a)—(c) is the content of [28, Theorem 4]. ∎
The class of C*-algebras with no infinite-dimensional irreducible representations has been investigated in [10]. Since every such a C*-algebra is necessarily liminary, we introduce the following:
Definition 2.3.
A separable C*-algebra is preliminary if it has no infinite-dimensional irreducible representation.
In the unital case, a liminary C*-algebra is necessarily preliminary [7, Section IV.1.3]. If $A$ is any preliminary nontrivial separable C*-algebra, then $A\otimes K\left(H\right)$ is liminary and not preliminary.
3. Some topology notions
In this section we recall some notions from topology, and introduce some new ones.
3.1. Fell-convergent nets
In a topological space that is not necessarily Hausdorff, a convergent sequence might have more than one limit. Furthermore, it might have accumulation points that are not points to which it converges. We consider the following terminology:
Definition 3.1.
Let $X$ be a topological space.
  • a sequence $\left(x_{n}\right)$ in $X$ is Fell-convergent if the (possibly empty) set of points to which converges is equal to the set of its accumulation points;
  • if $\left(x_{n}\right)$ is a Fell-convergent sequence, then we define its limit set to be the (possibly empty) set of points to which it converges.
Fell-convergent sequences are called primitive in [26, 25], and properly convergent in [4, 38] (as long as they have at least one limit).
3.2. Hausdorff decompositions
Recall that a topological space is Polish if it has a countable basis of open sets, and its topology is induced by a complete metric. The notion of Polish decomposition of a topological space has been considered by Dixmier, Blackadar, Lazar, and Taylor [37, 8, 17], under the name of Hausdorff decomposition.
Definition 3.2.
Let $X$ be a topological space. A Polish decomposition of $X$ of length an ordinal $\sigma$ is a chain $\left(U_{\beta}\right)_{\beta\leq\sigma}$ of open subspaces of $X$ such that:
  1. (1)
    for every $\alpha<\sigma$, $U_{\alpha+1}\setminus U_{\alpha}$ is nonempty and Polish with the subspace topology;
  2. (2)
    for every limit ordinal $\lambda\leq\sigma$, $U_{\lambda}$ is the union of $U_{\beta}$ for $\beta<\lambda$;
  3. (3)
    $U_{\sigma}=X$.
We introduce the notion of standard topological space. Recall the notion of standard Borel space from [34, Section 12.B].
Definition 3.3.
A topological space is standard if:
  1. (1)
    it has a countable basis of open sets, and
  2. (2)
    the Borel structure it generates is standard.
Lemma 3.4.
If $X$ is a standard topological space, then there is a Polish topology $\tau_{P}$ finer than the topology of $X$ that induces the same Borel structure.
Proof.
This follows from [34, Lemma 13.2 and Lemma 13.3] considering that $X$ has a countable basis of open sets. ∎
Lemma 3.5.
Let $X$ be a second countable topological space. If $\left(U_{\beta}\right)_{\beta\leq\sigma}$ is a Hausdorff decomposition of $X$, then $\sigma$ is a countable ordinal.
Proof.
This follows from [34, Theorem 6.9]. ∎
Proposition 3.6.
Let $X$ be a second countable topological space. The following assertions are equivalent:
  1. (1)
    $X$ has a Polish decomposition;
  2. (2)
    every nonempty closed subset $C$ of $X$ has a nonempty subset $U$ that is open in the subspace topology of $C$ and Polish with respect to the subspace topology.
    Furthermore, these assertions imply that $X$ is standard.
Proof.
[…]
$X_{\alpha}\cap C=\left(X_{\alpha}\setminus X_{\alpha-1}\right)\cap C$
is a nonempty. Furthermore, it is open (in the subspace topology of $C$) since $X_{\alpha}$ is open in $X$. It is also Polish (with respect to the induced subspace topology of $C$), being a closed subspace of the Polish space $X_{\alpha}\setminus X_{\alpha-1}$.
The fact that $X$ is standard is easily proved by induction on the length of a Polish decomposition. ∎
3.3. Constructible sets and Baire spaces
We consider the following terminology about collections of subsets of a given set; see [29, Section 4].
Definition 3.7.
Let $X$ be a set and $\mathcal{A}$ be a collection of subsets of $X$. Then $\mathcal{A}$ is :
  • a ring of sets if for $A,B\in\mathcal{A}$, $A\cap B\in\mathcal{A}$ and $A\setminus B\in\mathcal{A}$;
  • an algebra of sets if it is a ring of sets and closed under complements.
Let $X$ be a topological space. Recall that a locally closed set in $X$ is the intersection of an open set and a closed set. We will consider the more general class of constructible sets. By definition, these are the elements of the algebra of sets generated by the open sets [44].
Let $X$ be a topological space. A subspace $A$ of $X$ is called meager if it is contained in the union of a countable family of closed sets with empty interior. A topological space $X$ is Baire if every nonempty open subset $U$ of $X$ is not meager; see [34, Definition 8.2]. A topological space $X$ is hereditarily Baire if each of its closed subspace of $X$ is Baire, when endowed with the subspace topology. It is easily seen by induction on the complexity of $C$ that, if $X$ is a hereditarily Baire space, and $C\subseteq X$ is constructible, then $C$ is hereditarily Baire.
3.4. Polish presentations
We are interested in spaces, such as the primitive spectra of C*-algebras, that are not necessarily Polish. These spaces, however, satisfy a property that we isolate in the following:
Definition 3.8.
A space with a standard Polish presentation is a space $X$ explicitly presented as the quotient $P/E$ of a Polish space $P$ by an equivalence relation $E$ with the property that:
  1. (1)
    the quotient map $P\rightarrow P/E$ is (continuous and) open when $X$ is endowed with the quotient topology, i.e., the $E$-saturation of an open subset of $P$ is open;
  2. (2)
    the quotient Borel structure on $P/E$ induced by the Borel structure on $P$ is standard, and in particular $E$ is a smooth equivalence relation [34, Exercise 18.20].
A space with a standard Polish presentation has a countable basis of open sets, and it is hereditarily Baire. This follows from the fact that every Polish space is second countable and a Baire space [34, Theorem 8.4], and every $G_{\delta}$ subspace of a Polish space is Polish [34, Theorem 3.11].
Obviously, every Polish space has a Polish presentation. Thus, we can think of this notion as a generalization of the notion of Polish space, including spaces that are not necessarily Hausdorff.
3.5. Countably compactly based spaces
We consider the following property for topological spaces, which simultaneously subsumes second coutability and local compactness, and it is equivalent to their conjunction in the Hausdorff case. For a set $Y\subseteq X$ we denote by $\mathrm{int}\left(Y\right)$ its interior.
Definition 3.9.
Let $X$ be a (not necessarily Hausdorff) topological space. Let us say that $X$ countably compactly based if admits a countable collection $\mathcal{B}$ of compact sets such that for every open set $W$ in $X$ there exists a sequence $K_{n}$ in $\mathcal{B}$ such that
$W=\bigcup_{n\in\omega}\mathrm{int}\left(K_{n}\right)=\bigcup_{n\in\omega}K_{n}\text{.}$
We say that $X$ has a countable basis of compact open set when one can take the elements of $\mathcal{B}$ both compact and open.
It is easily seen by induction on the complexity that a constructible subspace of a countably compactly based space is countably compactly based.
4. The Fell compactification
In this section we recall a compactification construction due to Fell.
4.1. Fell spaces and their compactification
Let $X$ be a countably compactly based $T_{0}$ space. Denote by $F\left(X\right)$ the space of closed subsets of $X$. The Fell topology on $F\left(X\right)$, as defined by Fell in [26], is the topology that has as sub-basis of open sets the sets of the form
$\left\{C\in F\left(X\right):C\cap K=\varnothing\right\}$
and
$\left\{C\in F\left(X\right):C\cap U\neq\varnothing\right\}$
where $K\subseteq X$ is compact and $U$ is open; see also [17, 3.9.2]. In the Hausdorff case, this is also considered in [34, Exercise 12.7] and [6, Chapter 5].
It is proved in [26, Theorem 1] that $F\left(X\right)$ endowed with the Fell topology is a compact Hausdorff space, which is easily seen to be second countable when $X$ is countably compactly based. By identifying a point $x\in X$ with the closed subspace $\overline{\left\{x\right\}}$ of $X$, one can regard $X$ as a subspace of $F\left(X\right)$.
Definition 4.1.
Let $X$ be a Fell space. The Fell compactification $\Phi\left(X\right)$ is the closure of $X$ within $F\left(X\right)$.
It follows from the above remarks that $\Phi\left(X\right)$ is a compact second countable Hausdorff space. Its elements are characterized in [26, Lemma 2] as the limits sets of Fell-convergent sequences in $X$. Furthermore, a sequence $\left(x_{n}\right)$ in $X$ is Fell-convergent with limit set $C$ if and only if it converges to $C$ in $\Phi\left(X\right)$ [26, Lemma 2].
4.2. The Fell topology
Let $X$ be a countably compactly based $T_{0}$ space, which we regard as a subspace of its Fell compactification $\Phi\left(X\right)$.
Definition 4.2.
The Fell topology on $X$ is the subspace topology inherited from $\Phi\left(X\right)$, while the Jacobson topology on $X$ is its original topology.
Notice that the Fell topology on $X$ is in general finer than the Jacobson topology of $X$. Recall that a topological space is Polish if it has a countable basis of open sets and its topology is induced by a complete metric. A subspace of a Polish space is Polish with respect to the subspace topology if and only if it is $G_{\delta}$ [34, Theorem 3.11]. Every locally compact second countable Hausdorff space is Polish [34, Theorem 5.3].
Proposition 4.3.
Let $X$ be a countably compactly based $T_{0}$ space. Then $X$ is a $G_{\delta}$ subspace of $\Phi\left(X\right)$, whence a Polish space when endowed with the Fell topology. The Borel structure generated by the Fell topology on $X$ coincides with the Borel structure generated by the Jacobson topology. In fact, each open set in the Fell topology is a countable union of sets that are either closed or open in the Jacobson topology.
Proof.
Fix a compactible metric $d$ on $\Phi\left(X\right)$ with values in $\left[0,1\right]$. For $C\in F\left(X\right)\setminus\left\{\varnothing\right\}$, let $d_{C}$ be the function
$X\rightarrow\left[0,1\right]\text{, }x\mapsto d_{C}\left(x\right):=\inf_{c\in C}d\left(x,c\right)\text{.}$
Define $\mathrm{Lip}\left(X\right)$ to be the space of Lipschitz functions $X\rightarrow\left[0,1\right]$ with respect to $d$, excluding the function constantly equal to $1$. We consider $\mathrm{Lip}\left(X\right)$ as a (partially) ordered set with respect to the relation
$d\leq d^{\prime}\Leftrightarrow\forall x,y\in X\text{, }d\left(x,y\right)\leq d^{\prime}\left(x,y\right)\text{.}$
The assignment $C\mapsto d_{C}$ defines a function
$\Phi\left(X\right)\rightarrow\mathrm{Lip}\left(X\right)$
[…]
Observe now that, for $C\in\Phi\left(X\right)$, we have that
$C\in X\Leftrightarrow d_{C}\text{ is maximal in }Z\left(X\right)\text{.}$
Furthermore,
$R:=\left\{\left(D_{0},D_{1}\right)\in\Phi\left(X\right)\times\Phi\left(X\right):d_{D_{0}}\leq d_{D_{1}}\text{ or }d_{D_{1}}\leq d_{D_{0}}\right\}$
is compact. Hence, its complement
$S:=\left(\Phi\left(X\right)\times\Phi\left(X\right)\right)\setminus R$
is a countable union of compact sets. For $C\in\Phi\left(X\right)$ we have
$C\notin X\Leftrightarrow\exists\left(D_{0},D_{1}\right)\in S\text{ such that }C=\sup\left\{d_{D_{0}},d_{D_{1}}\right\}\text{.}$
This shows that $\Phi\left(X\right)\setminus X$ is a countable union of compact sets. Compact sets are closed in $\Phi\left(X\right)$, which is Hausdorff. Whence, $X$ is $G_{\delta}$ in $\Phi\left(X\right)$, and Polish when endowed with the Fell topology.
It remains to observe that, if $K\subseteq X$ is compact and $U\subseteq X$ is open with respect to the Jacobson topology, then
$\{x\in X:\overline{\left\{x\right\}}\cap U\neq\varnothing\}=U$
is open in the Jacobson topology. Furthermore, using the assumption that $X$ is countably compactly based,
$\{x\in X:\overline{\left\{x\right\}}\cap K=\varnothing\}$
is a countable union of sets that are closed in the Jacobson topology. ∎
Corollary 4.4.
If $X$ is a countably compactly based $T_{0}$ space, then $X$ is a standard topological space as in Definition 3.3.
Remark 4.5.
The proof above can be seen as a general version of the proof considered in [17, 1.9.13 and 3.9.2 and 3.9.3] in the case when $X=\mathrm{Prim}\left(A\right)$ for some separable C*-algebra AA; see also [19, Section 2].
Let $X$ be a countably compactly based $T_{0}$ space. A different topology on $F\left(X\right)$ was considered by Michael in [43]. This has a sub-basis of open sets consisting of sets of the form
$\left\{C\in F\left(X\right):C\cap K=\varnothing\right\}$
and
$\left\{C\in F\left(X\right):C\cap U\neq\varnothing\right\}$
as above, where $U$ is open and $K$ is closed (rather than compact) in $X$.
4.3. Universal property
A universal property of the Fell compactification was also isolated by Fell in [26, Theorem 2]. Let us say that a function $f:X\rightarrow Y$ between countably compactly based $T_{0}$ spaces is Fell-continuous if it maps a Fell-convergent sequence $\left(x_{n}\right)$ with limit $C$ to Fell-convergent sequence with limit $f\left(C\right)$. In the case of locally compact second countable Hausdorff spaces, this is equivalent to being proper.
Proposition 4.6.
Let $X$ and $Y$ be countably compactly based $T_{0}$ spaces, and $f:X\rightarrow Y$ be a function. The following assertions are equivalent:
  1. (1)
    $f$ is Fell-continuous;
  2. (2)
    $f$ is the restriction to $X$ to a continuous function $\Phi\left(X\right)\rightarrow\Phi\left(Y\right)$;
  3. (3)
    $f$ is continuous with respect to the Jacobson topologies, and $f^{-1}\left(K\right)$ is Jacobson-compact for every $K\subseteq Y$ Jacobson-compact.
Proof.
[…]
Suppose now that $K\subseteq Y$ is compact. We need to prove that $f^{-1}\left(K\right)$ is compact. Since $X$ is countably compactly based, it suffices to prove that $f^{-1}\left(K\right)$ is countably compact, i.e., if $\left(x_{n}\right)$ is a sequence in $f^{-1}\left(K\right)$, then $\left(x_{n}\right)$ has an accumulation point in $f^{-1}\left(K\right)$. After passing to a subsequence, we can assume that $\left(x_{n}\right)$ is Fell-convergent. Let $C$ be its limit set. Thus, $\left(f\left(x_{n}\right)\right)$ is Fell-convergent with limit set $f\left(C\right)$. Since $K$ is Jacobson-compact, and $\left(f\left(x_{n}\right)\right)$ is Fell-convergent, $f\left(C\right)\cap K\neq\varnothing$. Hence, $C\cap f^{-1}\left(K\right)\neq\varnothing$.
[…]
Corollary 4.7.
The category of compact Polish spaces is a reflective subcategory of the category of countably compactly based $T_{0}$ spaces and Fell-continuous functions. The left adjoint of the inclusion functor is the functor $X\mapsto\Phi\left(X\right)$ that assigns to a countably compactly based $T_{0}$ space its Fell compactification. The unit of the adjunction is the natural transformation $X\rightarrow\Phi\left(X\right)$ given by the inclusion.
4.4. One-point compactification
Let $X$ be a countably compactly based $T_{0}$ space. If $X$ is compact, define $X^{\dagger}=X$. If $X$ is not compact, define its one-point compactification $X^{\dagger}$ to be the space $X\cup\left\{\infty\right\}$ where $X$ is an open subspace and the neighborhoods of $\infty$ are the sets of the form
$\left(X\setminus K\right)\cup\left\{\infty\right\}$
where $K\subseteq X$ is compact. Then we have that $X^{\dagger}$ is a countably compactly based and compact $T_{0}$ space. Observe that $X^{\dagger}$ can be identified with a $G_{\delta}$ subspace of $\Phi\left(X\right)$, in such a way that $\infty$ corresponds to the empty set (which is indeed a limit set under the assumption that $X$ is not compact). By way of this identification, $\Phi(X^{\dagger})$ can be identified with $\Phi\left(X\right)$. Clearly, this definition of the one-point compactification coincides with the usual one in the Hausdorff case (in which case it also coincides with the Fell compactification, although they are different in general). The following universal property of the one-point compactification is an immediate consequence of the corresponding one for the Fell compactification; see Proposition 4.6.
Proposition 4.8.
Let $X$ and $Y$ be countably compactly based $T_{0}$ spaces, and $f:X\rightarrow Y$ be a function. The following assertions are equivalent:
  1. (1)
    $f$ is Fell-continuous;
  2. (2)
    the function $X^{\dagger}\rightarrow Y^{\dagger}$ mapping $\infty_{X}$ to $\infty_{Y}$ is Fell-continuous.
Corollary 4.9.
The category of compact countably based $T_{0}$ spaces and Fell-continuous functions is a reflective subcategory of the category of countably compactly based $T_{0}$ spaces and Fell-continuous functions. The left adjoint of the inclusion functor is the functor $X\mapsto X^{\dagger}$ that assigns to a countably compactly based $T_{0}$ space its one-point compactification. The unit of the adjunction is the natural transformation $X\rightarrow X^{\dagger}$ given by the inclusion.
4.5. Fell and Kuratowski
Let $X$ be a countably compactly based $T_{0}$ space. Let $\left(F_{n}\right)$ be a sequence of closed subsets of $X$. One defines
$\limsup_{n}F_{n}$
to be the set of $z\in X$ such that for every neighborhood $U$ of $z$ one has that, for infinitely many $n\in\mathbb{N}$, $F_{n}\cap U\neq\varnothing$. Likewise, one defines
$\liminf_{n}F_{n}$
to be the set of $z\in X$ such that for every neighborhood $U$ of $z$ one has that, for all but finitely many $n\in\mathbb{N}$, $F_{n}\cap U\neq\varnothing$. The sequence $\left(F_{n}\right)$ is Kuratowski-convergent to $F$ if
$\limsup_{n}F_{n}=\liminf_{n}F_{n}=F\text{.}$
Then for a countably compact $T_{0}$ space, a sequence $\left(F_{n}\right)$ of closed sets converges to $F$ in $F\left(X\right)$ with respect to the Fell topology if and only if it is Kuratowski–convergent to $F$; see [53, 45].
5. Preliminary spaces
In this section, we introduce the notion of weight function—or, briefly, weight—on a countably compactly based $T_{0}$ space. In terms of this notion, we define preliminary spaces.
5.1. Weight functions
We consider $\overline{\mathbb{N}}=\mathbb{N}\cup\left\{\infty\right\}$ to be a topological space, with the topology that has the sets $n,\infty]$ as open sets, which we call the Jacobson topology. Then $\overline{\mathbb{N}}$ is a countably compact $T_{0}$ space.
Definition 5.1.
Let $w:X\rightarrow\overline{\mathbb{N}}$ be a function. Then $w$ is:
  • a weight if the extension $w:\Phi\left(X\right)\rightarrow\overline{\mathbb{N}}$ defined by setting
    $w\left(C\right):=\sum_{y\in C}w\left(y\right)$
    for every limit set $C\in\Phi\left(X\right)$ is continuous with respect to the Fell topology on $\Phi\left(X\right)$ and the Jacobson topology on $\mathbb{N}$;
  • finite if $w\left(x\right)\in\mathbb{N}$ for every $x\in X$.
Remark 5.2.
In Definition 5.1, we adopt the convention that an empty sum is zero, so that $w\left(\infty\right)=0$.
In the following lemma, we prove that the condition of being a weight is equivalent to a seemingly weaker condition, which can be formulated solely in terms of Fell-convergent sequences in $X$.
Lemma 5.3.
Suppose that $w:X\rightarrow\overline{\mathbb{N}}$ is a function. The following assertions are equivalent:
  1. (1)
    $w$ is a weight;
  2. (2)
    for every Fell-convergent sequence $\left(x_{n}\right)$ in $X$ with limit set $C$,
    $\sum_{y\in C}w\left(y\right)\leq\liminf_{n}w\left(x_{n}\right)\text{.}$
Proof.
[…]
Suppose that $\left(C_{n}\right)$ is a sequence in $\Phi\left(X\right)$ converging to $C$ and such that $\phi\left(C_{n}\right)=d\in\mathbb{N}$ for every $n\in\mathbb{N}$. Fix $\ell\geq 1$ and $y_{1},\ldots,y_{\ell}\in C$. Then there exist open neighborhoods $U_{1},\ldots,U_{\ell}$ of $y_{1},\ldots,y_{\ell}$, respectively, such that
$w|_{U_{i}}\geq w\left(y_{i}\right)$
for $1\leq i\leq\ell$. Considering the open neighborhood
$\left\{D\in\Phi\left(C\right):U_{i}\cap D\neq\varnothing\right\}$
of $C$, we observe that we can assume that there exist $y_{n}^{\left(i\right)}\in U_{i}\cap C_{n}$ for all $n\in\mathbb{N}$ and $i\in\left\{1,2,\ldots,\ell\right\}$. After passing to a subsequence, we can assume that there is a partition $A_{1},\ldots,A_{m}$ of $\left\{1,\ldots,\ell\right\}$ and, for each $1\leq k\leq m$, $i_{k}\in\left\{1,2,\ldots,\ell\right\}$ such that:
  1. (1)
    $(y_{n}^{\left(i_{k}\right)})$ has all the elements of $A_{k}$ as limits;
  2. (2)
    $y_{n}^{i_{1}},\ldots,y_{n}^{i_{m}}$ are pairwise distinct.
After passing to subsequences, we can assume that $(y_{n}^{\left(k_{i}\right)})$ is Fell-convergent. By hypothesis, we have
$\sum_{z\in A_{k}}w\left(z\right)\leq\liminf_{n}w(y_{n}^{(k_{i})})$
and hence
$\sum_{i=1}^{\ell}w\left(y_{i}\right)$ $\leq$ $\sum_{k=1}^{m}\sum_{z\in A_{k}}w\left(z\right)\leq\sum_{k=1}^{m}\liminf_{n}w(y_{n}^{(i_{k})})$
$\leq$ $\liminf_{n}\sum_{k=1}^{m}w(y_{n}^{\left(i_{k}\right)})\leq\liminf_{n}w\left(C_{n}\right)\leq d\text{.}$
Since this holds for every $\ell\in\mathbb{N}$ and $y_{1},\ldots,y_{\ell}\in C$ we conclude that $w\left(C\right)\leq d$. ∎
5.2. Preliminary spaces
We now introduce the notion of preliminary space in terms of the notions above.
Definition 5.4.
A weighted space is a topological space endowed with a distinguished weight. A preliminary space is a weighted space $\left(X,w\right)$ such that:
  1. (1)
    $X$ is countably compactly based and $T_{0}$;
  2. (2)
    $X$ is hereditarily Baire;
  3. (3)
    $w$ is finite.
The preliminary space $\left(X,w\right)$ is $n$-homogeneous when $w$ is the rank function constantly equal to $n$.
Remark 5.5.
If $\left(X,w\right)$ is a liminary space, and $C\subseteq X$ is constructible, then again it is established by induction on the complexity of $C$ that $C$ is a preliminary space when endowed with the induced topology and the restriction of $w$ to $C$.
Remark 5.6.
An $n$-homogeneous preliminary space must necessarily be Polish.
As a preliminary space is, by definition, countably compactly based and $T_{0}$, it is also a standard topological space by Corollary 4.4. We now show that it also admits a Polish decomposition in the sense of Definition 3.2.
Lemma 5.7.
Let $\left(X,w\right)$ be a preliminary space. Then $X$ admits a Polish decomposition.
Proof.
By Lemma 3.6, it suffices to prove that every closed subspace of $X$ admits a Polish open subset. Since a closed subspace of $X$ is still preliminary, it suffices to prove that $X$ has a Polish open subset. For $n\in\mathbb{N}$ define
$X_{n}:=\left\{x\in X:w\left(x\right)\leq n\right\}\text{.}$
Then we have that $X$ is the union of $X_{n}$ for $n\in\mathbb{N}$. Since by hypothesis $X$ is a Baire space, there exists $n\in\omega$ such that $X_{n}$ contains a nonempty open subset $U$. Let
$k:=\min\left\{i\in\mathbb{N}:U\cap X_{i}\neq\varnothing\right\}\text{.}$
Then we have that
$V:=U\cap X_{k}=U\cap\left(X_{k}\setminus X_{k-1}\right)$
is a nonempty open subset $X$. Furthermore, $w|_{V}$ is constantly equal to $k$. Thus, $V$ is a $k$-homogeneous preliminary space, and hence Polish. ∎
5.3. Cantor–Bendixon analysis of preliminary spaces
Suppose that $\left(X,w\right)$ is a preliminary space. We define a canonical decomposition of $X$, which can be seen as a sort of Cantor–Bendixon analysis of $X$ [34, Theorem 6.4].
We say that an element of $X$ has order 0 and rank $n<\omega$ if it has a neighborhood $U$ that is $n$-homogeneous, i.e., $w|_{U}$ is constantly equal to $n\in\mathbb{N}$. Define then $X_{n}$ to be the (open) set of elements of order 0 and rank $n$. The derivative $X^{1}$ is the (closed) complement of the union of $X_{n}$ for $n\in\mathbb{N}$.
Recursively, one sets
$X^{\alpha+1}:=\left(X^{\alpha}\right)^{1}\text{.}$
For $\lambda$ limit one sets
$X^{\lambda}=\bigcap_{\beta<\alpha}X^{\beta}\text{.}$
By [34, Theorem 6.9] there exists $\sigma<\omega_{1}$ such that $X^{\sigma}=\varnothing$. The least such a $\sigma$ is called the order of $X$. When $X$ is compact, $\sigma$ must be a successor ordinal. The set $X^{\alpha}$ is called the $\alpha$-th derivative of $X$. An element $x\in X$ has order $\alpha$ and rank $\omega\alpha+n$ if it belongs to $X_{n}^{\alpha}$, i.e., there exists a neighborhood $U$ of $x$ such that $U\cap X^{\alpha}$ is $n$-homogeneous. The rank of XX\ is the supremum of the ranks of its elements.
The proof of the following lemma is easily established by induction.
Lemma 5.8.
Let $X,Y$ be preliminary spaces. Let $f:X\rightarrow Y$ be a continuous function such that $w_{X}\leq w_{Y}\circ f$. Then for every $x\in X$,
$r_{w_{X}}\left(x\right)\leq r_{w_{Y}}\left(f\left(x\right)\right)$
and in particular
$r\left(X\right)\leq r\left(Y\right)\text{.}$
5.4. Cantor–Bendixon analysis of scattered spaces
The Cantor–Bendixon analysis of a scattered locally compact Polish space [34, Section 6.C] can be seen as a particular instance of the one described above. Indeed, let $X$ be a scattered (equivalently, countable) locally compact Polish space. Recall that one defines its derivative
$X^{1}=\left\{x\in X:x\text{ is not an isolated point}\right\}\text{.}$
Recursively, on sets
$X^{\alpha+1}:=\left(X^{\alpha}\right)^{1}$
and
$X^{\lambda}=\bigcap_{\beta<\lambda}X^{\beta}\text{.}$
One can consider $X$ as a weighted space as follows. For each $\alpha<\omega_{1}$, fix an enumeration $\left(x_{n}^{\alpha}\right)$ of $X^{\alpha+1}\setminus X^{\alpha}$, where $X^{0}=\varnothing$. One can then define
$w\left(x_{n}^{\alpha}\right):=n\text{.}$
This turns $X$ into a preliminary space, in such a way that its derivatives coincide with the Cantor–Bendixon derivatives.
6. Spectra of C*-algebras
In this section we relate the primitive spectra of C*-algebras with the topological notions introduced above.
6.1. C*-algebra spectra
We record in the following proposition some properties of spectra of C*-algebras. Suppose that $A$ is a separable C*-algebra. Let $P(A)$ be the pure state space of $A$ [48, Section 4.3]. By [48, Proposition 4.3.2], $P\left(A\right)$ is a Polish space when endowed with the w*-topology. We have a canonical map
$P\left(A\right)\rightarrow X\text{, }\phi\mapsto\mathrm{\mathrm{Ker}}\left(\pi_{\phi}\right)$
mapping a pure state to the kernel of the corresponding irreducible representation.
Proposition 6.1.
Let $A$ be a separable C*-algebra, and $X:=\mathrm{Prim}\left(A\right)$:
  1. (1)
    $X$ is a countably compactly based and hereditarily Baire $T_{0}$ space;
  2. (2)
    The map $P\left(A\right)\rightarrow X$, $\phi\mapsto\mathrm{\mathrm{Ker}}\left(\pi_{\phi}\right)$ is a Polish presentation.
Proof.
(1) If $I$ and $J$ are distinct primitive ideals of $A$, then either $I$ is not contained in $J$ or vice versa. This shows that $X$ is $T_{0}$.
Let $\mathcal{A}$ be a countable dense subset of $\mathrm{\mathrm{Ball}}(A_{+})$. For $x\in A_{+}$ define
$U\left(x\right):=\left\{J:\check{x}\left(J\right)>1/2\right\}$
and
$K\left(x\right):=\left\{J:\check{x}\left(J\right)\geq 1/2\right\}$
Then we have that $U\left(x\right)$ is open and $K\left(x\right)$ is compact [7, Proposition II.6.5.6]. Let $\mathcal{B}$ be the collection of compact sets $K\left(x\right)$ for $x\in\mathcal{A}$. The proof of [7, Proposition II.6.5.6(ii)] shows that $\mathcal{B}$ witnesses that $X$ is countably compactly based.
(2) This map is continuous and open by [48, Theorem 4.3.3]. Furthermore, by [17, 3.9.2], the corresponding quotient Borel structure on $P\left(A\right)$ is standard, and coincides with the Borel structure generated by the (Polish) Fell topology on $\mathrm{Prim}\left(A\right)$. ∎
6.2. The Fell compactification of the primitive spectrum
Let $A$ be a separable C*-algebra. Consider the countably compactly based $T_{0}$ space $X:=\mathrm{Prim}\left(A\right)$ endowed with the Jacobson topology. In this case, we can identify the space $F\left(X\right)$ of closed subspaces of $X$ with the space of all closed ideals of $A$. In turn, the space of closed ideals of $A$ can be identified with the space of C*-seminorms on $A$, by mapping a closed ideal $I$ to the C*-seminorm
$N_{I}:A\rightarrow\mathbb{R}_{+}\text{, }a\mapsto\left\|a+I\right\|_{A/I}\text{;}$
see [17, 1.9.13]. With respect to this correspondence, the Fell topology on $F\left(X\right)$ corresponds to the weakest topology that renders all the functions
$a\mapsto N_{I}\left(a\right)$
continuous for $a\in A$, which is called the strong topology in [3]. The Michael topology on $F\left(X\right)$ corresponds to the weakest topology that makes the functions $a\mapsto N_{I}\left(a\right)$ lower semi-continuous, called the weak topology in [3]. Equivalently, it is the topology that has the sets $F\left(X\right)\setminus\mathrm{hull}\left(J\right)$, where $J$ ranges among the closed ideals of $A$, as sub-basis of closed sets. In particular, the corresponding subspace topology on $X$ is the Jacobson topology.
The (standard) Borel structure induced by the Jacobson or, equivalently, Fell topology on $X$ coincides with the quotient Borel structure induced from the Borel structure on $P\left(A\right)$ via the surjective map $P\left(A\right)\rightarrow X$ as in Proposition 6.1(2).
6.3. Unitization
Let $A$ be a nonunital C*-algebra. Its unitization $A^{\dagger}$ [7, Section II.1.2] is defined to be the algebra $A\oplus\mathbb{C}$ with coordinate-wise addition and scalar multiplication, and multiplication induced by the identification
$\left(a,\lambda\right)=a+\lambda 1$
for $a\in A$ and $\lambda\in\mathbb{C}$, where $1$ acts as an identity. The C*-norm is defined to be the norm of elements of $A^{\dagger}$ as left multiplication operators on $A$. Equivalently, fixing a faithful nondegenerate representation $\subseteq B\left(H\right), AA^{\dagger}$ is the C*-subalgebra of $B\left(H\right)$ generated by $A$ and $1$. When $A$ is unital, one sets $A^{\dagger}=A$.
This construction can be seen as a noncommutative analogue of the one-point compactification of a space. Indeed, if $A=C_{0}\left(X\right)$ for some locally compact non compact Polish space, then $A^{\dagger}\cong C\left(X^{\dagger}\right)$ where $X^{\dagger}$ is the one-point compactification of $X$; see [7, II.1.2].
Recall that an element $a$ of a C*-algebra $A$ is:
  • full if the closed ideal generated by $a$ is all of $A$ [7, II.5.3.10];
  • strictly full if $\left(a-\varepsilon\right)_{+}$ is full for some (hence, for all sufficiently small) $\varepsilon>0$ [46].
If $A$ is a separable C*-algebra, then $\mathrm{Prim}\left(A\right)$ is compact if and only if $A$ has a strictly full element; see [47, Lemma 2.1].
Proposition 6.2.
Let $A$ be a separable C*-algebra with primitive spectrum $X$. Suppose that $X$ is not compact. Then the function $\phi:X^{\dagger}\rightarrow\mathrm{Prim}\left(A^{\dagger}\right)$ mapping $\infty$ to $A=\mathrm{Ker}\left(\chi\right)$, where $\chi$ is the character $a+\lambda 1\mapsto\lambda$ is a homeomorphism.
Proof.
Considering the extension
$A\rightarrow A^{\dagger}\rightarrow\mathbb{C}$
shows that $\phi$ is a bijection. Furthermore, $\phi|_{X}:X\rightarrow\mathrm{Prim}\left(A\right)$ is a homeomorphism, and in particular Fell-continuous.
Suppose that $\left(t_{n}\right)$ is a sequence in $X$. Set $J_{n}:=\phi\left(t_{n}\right)\in\mathrm{Prim}\left(A\right)$. Then the universal property of the one-point compactification shows that, if $\left(t_{n}\right)$ converges to $\infty$, then $\left(J_{n}\right)$ converges to $\phi\left(\infty\right)=A$.
Conversely, suppose that $\left(J_{n}\right)$ converges to $A=\mathrm{\mathrm{Ker}}\left(\chi\right)$. We need to prove that $\left(t_{n}\right)$ has no accumulation point in $\mathrm{Prim}\left(A\right)$. After passing to a subsequence, it suffices to prove that its limit set in $X$ is empty. Suppose that $t_{n}$ converges to $t\in X$. Then we have that $J_{n}$ converges to $J=\phi\left(t\right)$. This means that
$\left\|x/J\right\|_{A/J}\leq\liminf_{n}\left\|x/J_{n}\right\|$
for every $x\in A$. Observe that the sets
$\mathrm{\mathrm{Ker}}\left(\chi\right)\cup\left\{J\in\mathrm{Prim}\left(A\right):\left\|a/J\right\|<\lambda\right\}$
for $a\in A_{+}$ and $\lambda>0$ are neighborhoods of $\mathrm{\mathrm{Ker}}\left(\chi\right)$ in $\mathrm{Prim}\left(A^{\dagger}\right)$. Since $\left(J_{n}\right)$ converges to $\mathrm{\mathrm{Ker}}\left(\chi\right)$, this implies that, for $x\in A_{+}$,
$\left\|x/J\right\|\leq\liminf_{n}\left\|x/J_{n}\right\|=0\text{.}$
This shows that $J=A$ contradicting the assumption that $J=\phi\left(t\right)\in\mathrm{Prim}\left(A\right)$. ∎
Let $A$ be a preliminary separable C*-algebra. Then it is easily seen that its unitization $A^{\dagger}$ is still preliminary. Indeed, $A$ can be realized as a C*-subalgebra of a product of matrix algebras. Thus, the same holds for its unitization, whence $A^{\dagger}$ is preliminary as well. By ways of this observation, one can reduce some problems about preliminary C*-algebras to the unital case.
6.4. Spectra of preliminary C*-algebras
We show that the spectrum of a preliminary C*-algebra is indeed a preliminary space.
Proposition 6.3.
Let $A$ be a separable preliminary C*-algebra with primitive spectrum $X$. Define the function
$w:X\rightarrow\mathbb{N}\text{, }\mathrm{\mathrm{Ker}}\left(\pi\right)\mapsto\mathrm{\dim}\left(\pi\right)\text{.}$
Then $w$ is a finite weight on $X$, and $\left(X,w\right)$ is a preliminary space.
Proof.
By Proposition 6.2, after replacing $A$ with its unitization, we can assume that it is unital. In this case, we have that
$\dim\left(\pi\right)=\mathrm{Tr}\left(\pi\left(1_{A}\right)\right)$
for every irreducible representation $\pi$ of $A$. Thus, it follows from [4, Theorem 2.4] that $w$ is a weight, which is finite by definition of preliminary C*-algebra. The conclusion now follows from Proposition 6.1. ∎
We consider the spectrum of a preliminary C*-algebra as a preliminary space endowed with the canonical weight function defined in Proposition 6.3.
7. The Fell and Pedersen rank
In this section we define a notion of Fell rank for representations of a separable C*-algebra, which we use to define the Fell rank the C*-algebra itself. In turns, this yields a notion of type $I_{\alpha}$ separable C*-algebra for an arbitrary countable ordinal $\alpha$. We also define the Pedersen rank of positive elements and of C*-algebras, which in turn yields the notion of $\alpha$-subhomogeneous C*-algebra.
7.1. The Pedersen weight
We consider a canonical weight associated with elements of a C*-algebra.
Definition 7.1.
Let $A$ be a separable C*-algebra with primitive spectrum $X$. The Pedersen weight corresponding to $a\in A$ is the function
$w_{a}:X\rightarrow\overline{\mathbb{N}}\text{, }\mathrm{\mathrm{Ker}}\left(\rho\right)\mapsto\mathrm{\mathrm{\mathrm{rank}}}\left(\rho\left(a\right)\right)\text{.}$
The Pedersen weight of $A$ is the function
$w_{A}:X\rightarrow\overline{\mathbb{N}}\text{, }\mathrm{\mathrm{Ker}}\left(\rho\right)\mapsto\mathrm{\dim}\left(\rho\right)\text{.}$
Remark 7.2.
It is obvious that $w_{a}\leq w_{A}$ for every $a\in A$.
In the following proposition we verify that the Pedersen weight is indeed a weight.
Proposition 7.3.
Let $A$ be a separable C*-algebra, and $a\in A$.
  1. (1)
    $w_{A}$ is a weight;
  2. (2)
    $w_{a}$ is a weight;
  3. (3)
    AA is preliminary if and only if $w_{A}$ is finite.
Proof.
(1) follows from [4, Corollary 2.5].
(2) Suppose that $\left(\mathrm{\mathrm{Ker}}(\rho_{n})\right)$ is a Fell-convergent sequence with limit set $C$. We need to show that
$\sum_{J\in C}w_{b}\left(J\right)\leq\liminf_{n}w\left(\mathrm{ker}\left(\rho_{n}\right)\right)=\liminf_{n}\mathrm{\mathrm{rank}}\left(\rho_{n}\left(b\right)\right)\text{.}$
We can assume that
$\mathrm{\mathrm{rank}}(\rho_{n}\left(b\right))=d\in\mathbb{N}$
for all $n\in\mathbb{N}$. Let $h\left(b,A\right)=\overline{b^{\ast}Ab}$ be the hereditary C*-subalgebra of $A$ generated by $b$. Recall the correspondence between irreducible representations of $h\left(b,A\right)$ and irreducible representations of $A$ that do not vanish on $h\left(b,A\right)$ [48, Section 4.1]. By this correspondence, without loss of generality, one can replace $A$ with $h\left(b,A\right)$. After doing so, one obtains that $\mathrm{\dim}\left(\rho_{n}\right)=d$ for every $n\in\mathbb{N}$. In turn, this implies by [4, Corollary 2.5] that $C=\left\{\mathrm{\mathrm{Ker}}\left(\pi_{1}\right),\ldots,\mathrm{\mathrm{Ker}}\left(\pi_{\ell}\right)\right\}$ is a finite set with at most $n$ elements. Furthermore,
$\sum_{i=1}^{\ell}\mathrm{\dim}\left(\pi_{\ell}\right)\leq n\text{.}$
Therefore,
$\sum_{J\in C}w_{b}\left(J\right)=\sum_{i=1}^{\ell}\mathrm{\mathrm{rank}}\left(\pi_{i}\left(b\right)\right)\leq\sum_{i=1}^{\ell}\mathrm{\dim}\left(\pi_{i}\right)\leq n\text{.}$
This concludes the proof.
(3) is definitional. ∎
7.2. Fell’s condition of order $\infty$
Let $A$ be a separable C*-algebra with primitive spectrum $X$, and $\pi_{0}$ be an irreducible representation of $A$. Let us say that a local section for $A$ is a pair $\left(b,U\right)$ where $b\in A$, $U\subseteq X$ is open, and $\rho\left(b\right)\neq 0$ for every $\mathrm{\mathrm{Ker}}\left(\rho\right)\in U$. This is a local section around $\pi_{0}$ if $\mathrm{\mathrm{Ker}}(\pi_{0})\in U$. Recall the definition of the Pedersen weigh $w_{b}$ associated with an element $b$ of $A$ as in Definition 7.1.
Definition 7.4.
Let $A$ be a separable C*-algebra, and $\pi_{0}$ an irreducible representation of $A$. Then $\pi_{0}$ satisfies Fell’s condition of order $\infty$ if there exists a local section $\left(b,U\right)$ around $\pi_{0}$ such that
$w_{b}|_{U}:U\rightarrow\overline{\mathbb{N}}\text{, }\mathrm{\mathrm{Ker}}\left(\rho\right)\mapsto\mathrm{\mathrm{rank}}\left(\rho\left(b\right)\right)$
is finite.
It turns out that such a condition is automatically satisfied whenever $\pi_{0}$ has a neighborhood of CCR representations. This is essentially established in the proof of [24, Lemma 2.5].
Proposition 7.5.
Let $A$ be a separable C*-algebra with primitive spectrum $X$, and $\pi_{0}$ an irreducible representation of $A$. Suppose that $\mathrm{\mathrm{Ker}}(\pi_{0})$ has a neighborhood $W$ in $X$ such that every irreducible representation $\rho$ of $A$ with $\mathrm{\mathrm{Ker}}\left(\rho\right)\in W$ is CCR. Then there exists a positive element $z\in A$ and a neighborhood $U$ of $\mathrm{\mathrm{Ker}}\left(\pi_{0}\right)$ contained in $W$ such that:
  1. (1)
    $\pi_{0}\left(z\right)$ is a rank-one projection, and
  2. (2)
    the weight $w_{z}|_{U}$ is finite.
In particular, $\pi_{0}$ satisfies Fell’s condition of order $\infty$.
Proof.
Let $P\in K\left(H_{\pi_{0}}\right)$ be a rank-one projection. Pick a positive element $z\in A$ such that $\pi_{0}\left(z\right)=P$. Fix $n\geq 2$ and set $\varepsilon:=1/n$. Let $F$ be a continuous non-negative function on $\mathbb{R}$ which satisfies
$F|_{\left(-\infty,1-\varepsilon\right)\cup\left(1+\varepsilon,+\infty\right)}\equiv 0$
and
$F\left(1\right)=1\text{.}$
Then we have that
$\pi_{0}\left(F\left(z\right)\right)=F(\pi_{0}\left(z\right))=F\left(P\right)=P\text{.}$
Considering the open condition
$\left\|F\left(z\right)+J\right\|>1-\varepsilon$
we see that there exists a neighborhood $U$ of $\mathrm{\mathrm{Ker}}\left(\pi_{0}\right)$ contained in $W$ such that
$\left\|F\left(z\right)+J\right\|>1-\varepsilon$
and in particular
$F\left(z\right)+J\neq 0$
for all $J\in U$. By assumption, every irreducible representation $\rho$ with $\mathrm{\mathrm{Ker}}\left(\rho\right)\in W$ is CCR. Thus, for $J\in U$ we have that $F\left(z\right)+J$ is a compact positive operator with all nonzero eigenvalues in $1-\varepsilon,1+\varepsilon]$. This implies that $F\left(z\right)+J$ has finite rank.
This shows that $\left(z,U\right)$ is a local section around $\pi_{0}$ such that $w_{z}|_{U}$ is finite, and it is a weight by Proposition 7.5. ∎
Recall that an irreducible representation $\pi_{0}$ of a separable C*-algebra satisfies Fell’s condition of order $1$ if there exists a local section $\left(b,U\right)$ around $\pi_{0}$ such that the weight $w_{b}|_{U}$ is finite of rank at most $1$. More generally, $\pi_{0}$ satisfies Fell’s condition of order $k<\omega$ as in [21, Section 3.2] if there exists a local section $\left(b,U\right)$ around $\pi_{0}$ such that the weight $w_{b}|_{U}$ is finite of rank at most $k$.
7.3. The Fell rank
Let $A$ be a separable C*-algebra. We define the Fell rank of $A$ and of its irreducible representations in terms of the Fell weight function.
Definition 7.6.
Let $A$ be a separable C*-algebra.
  • For a local section $\left(b,U\right)$, define the Fell rank $r_{F}\left(b,U\right)$ to be the rank of the weight $w_{b}|_{U}$, if such a weight is finite, and $\infty$ otherwise;
  • For an irreducible representation $\pi$ of $A$, the Fell rank $r_{F}(\pi)$ is the least of the Fell ranks of the local sections of $A$ around $\pi$;
  • The Fell rank $r_{F}\left(A\right)$ of $A$ is the supremum of the Fell ranks of its irreducible representations.
7.4. Pedersen order of positive elements
Recall that an element $x$ of a separable C*-algebra $A$ is abelian if and only if the hereditary C*-algebra $\overline{x^{\ast}Ax}$ is commutative. This is equivalent to the assertion that $\mathrm{\mathrm{rank}}\left(\pi\left(x\right)\right)\leq 1$ for every irreducible representation $\pi$ of $A$, i.e., the Pedersen weight $w_{x}$ is finite of rank at most $1$. More generally, $x$ has global rank at most $k<\omega$ as in [21, Section 3.2] if $w_{x}$ is finite of rank at most $k$. We consider the following generalization of these notions:
Definition 7.7.
Let $A$ be a separable C*-algebra with primitive spectrum $X$.
  • the Pedersen rank $r_{P}\left(x\right)$ of $x$ is the rank of the weight $w_{x}$, if such a weight is finite, and $\infty$ otherwise;
  • the Pedersen rank $r_{P}\left(A\right)$ is the supremum of the Pedersen ranks of the elements of $A$.
Remark 7.8.
Clearly, when $A$ is unital, its Pedersen rank is equal to the Pedersen rank of $1_{A}$.
Remark 7.9.
If $B$ is a hereditary C*-subalgebra of $A$, then every irreducible representation of $B$ is unitarily equivalent, for some irreducible representation $\pi$ of $A$ whose kernel does not contain $B$ [48, Proposition 4.1.9], to the restriction of $\pi|_{B}$ to its essential subspace [7, II.6.1.5]. Therefore, if $b\in B$ then its Pedersen rank is the same when computed in $B$ or in $A$.
Remark 7.10.
Considering that, for $b,c\in A$, and irreducible representation $\pi$,
$\mathrm{\mathrm{rank}}\left(\pi\left(bc\right)\right)\leq\min\left\{\mathrm{\mathrm{rank}}\left(\pi\left(b\right)\right),\mathrm{rank}\left(\pi\left(c\right)\right)\right\}$
we see that
$w_{bc}\leq\min\left\{w_{b},w_{c}\right\}\text{.}$
Thus
$r_{P}\left(bc\right)\leq\min\left\{r_{P}\left(b\right),r_{P}\left(c\right)\right\}\text{.}$
7.5. Fell ideals
Let $A$ be a separable C*-algebra, and $\alpha<\omega_{1}$. We define $P_{\alpha}\left(A\right)$ to be the set of elements of $A$ of Pedersen rank at most $\alpha$. Notice that $P_{\alpha}\left(A\right)$ is closed under products, so that the C*-subalgebra generated by $P_{\alpha}\left(A\right)$ is equal to the closed subspace generated by $P_{\alpha}\left(A\right)$, and it is in fact an ideal.
Definition 7.11.
Let $A$ be a separable C*-algebra, and $\alpha$ be a countable ordinal. We define the Fell ideal $\Pi_{\alpha}\left(A\right)$ of $A$ of rank $\alpha$ to be the closed ideal generated by $P_{\alpha}\left(A\right)$.
Lemma 7.12.
Let $A$ be a separable C*-algebra, and $\alpha<\omega_{1}$. Then $\Pi_{\alpha}\left(A\right)$ is liminary.
Proof.
If $\pi$ is an irreducible representation of $\Pi_{\alpha}\left(A\right)$, then it extends to an irreducible representation of $A$. Thus, if $a\in P_{\alpha}\left(A\right)$, then $\pi\left(a\right)$ is a finite-rank operator. It follows that if $a\in\Pi_{\alpha}\left(A\right)$, then $\pi\left(a\right)$ is a compact operator. ∎
7.6. Type $I_{\alpha}$ C*-algebras and $\alpha$-subhomogeneous C*-algebras
The following definition subsumes the notion of type $I_{n}$ C*-algebra for $n<\omega$ from [21, Section 3.2], which in turn has the notion of type $I_{0}$ C*-algebra as particular case. It also subsumes the notion of $n$-subhomogeneous C*-algebra, which is recovered as a particular case when $\alpha=n<\omega$.
Definition 7.13.
Let $A$ be a separable C*-algebra, and let $\alpha$ be a countable ordinal.
  • AA is type $I_{\alpha}$ if and only if $A=\Pi_{1+\alpha}\left(A\right)$;
  • AA is $\alpha$-subhomogeneous if and only if the Pedersen rank of $A$ is at most $\alpha$.
8. Fell and Pedersen
In this section, we characterize type $I_{\alpha}$ C*-algebras in terms of the Fell rank of elements, generalizing [21, Proposition 17].
8.1. Fell rank and Pedersen rank
As a preliminary observation, we describe the Fell rank of an irreducible representation in terms of elements of a given Pedersen rank. In the case when $\alpha=n<\omega$ we recover [21, Lemma 16]. Indeed, we follow a similar argument.
Proposition 8.1.
Let $A$ be a separable C*-algebra with primitive spectrum $X$, $\pi$ an irreducible representation of $A$, and $\alpha<\omega_{1}$. The following assertions are equivalent:
  1. (1)
    the Fell rank of $\pi$ is at most $\alpha$;
  2. (2)
    there exists $a\in A$ of Pedersen rank at most $\alpha$ such that $\pi\left(a\right)\neq 0$.
Proof.
[…]
$w_{a}\leq w_{b}$
and hence, considering the corresponding ranks,
$r_{P}\left(a\right)=r(w_{a})\leq r\left(w_{b}\right)\leq\alpha\text{.}$
[…]
8.2. Characterization of Fell ideals
Using the characterization of the Fell rank in terms of elements of given Pedersen rank, we can characterize the Fell ideals as follows.
Proposition 8.2.
Suppose that $A$ is a separable C*-algebra with primitive spectrum $X$, and $\alpha<\omega_{1}$. Let $\Pi_{\alpha}\left(A\right)$ be the Fell ideal of rank $\alpha$ of $A$. Then the closed subset $\mathrm{hull}\left(\Pi_{\alpha}\left(A\right)\right)$ of $X$ comprises the kernels of irreducible representations of Fell order greater than $\alpha$. In other words, for every irreducible representation $\pi$ of $A$,
$r_{F}\left(\pi\right)>\alpha\Leftrightarrow\Pi_{\alpha}\left(A\right)\subseteq\mathrm{\mathrm{Ker}}\left(\pi\right)\text{.}$
In particular, $\Pi_{\alpha}\left(A\right)$ is the largest ideal of $A$ of Fell order at most $\alpha$.
Proof.
This is an immediate consequence of Proposition 8.1. ∎
As a consequence of Proposition 8.2 we can characterize type $I_{\alpha}$ C*-algebras in terms of their Fell rank.
Theorem 8.3.
Let $A$ be a separable C*-algebra. The following assertions are equivalent:
  1. (1)
    AA is type $I_{\alpha}$;
  2. (2)
    AA has Fell rank at most $1+\alpha$.
Proof.
This is an immediate consequence of Proposition 8.2. ∎
8.3. Pedersen order and hereditary subalgebras
Let $A$ be a separable C*-algebra, and $x\in A$. The hereditary C*-subalgebra generated by $x$ is $\overline{x^{\ast}Ax}$ [7, Proposition II.3.4.2]. The following characterization of the Pedersen rank of an element generalizes [7, IV.1.1.7], which can be seen as the particular instance in the case $\alpha=1$, and [21, Lemma 16], which can be seen as the particular case when $\alpha=n<\omega$.
Proposition 8.4.
Let $A$ be a separable C*-algebra and $x\in A$. Define $B$ to be the hereditary C*-subalgebra of $A$ generated by $x$. Then the Pedersen rank of $x$ is equal to the Pedersen rank of $B$.
Proof.
Let $X$ be the primitive spectrum of $A$, and $Y$ the primitive spectrum of $B$. Consider the continuous function
$Y\rightarrow X\setminus\mathrm{hull}\left(B\right)\text{, }\mathrm{\mathrm{Ker}}\left(\pi\right)\mapsto\mathrm{\mathrm{Ker}}\left(\rho_{\pi}\right)$
from [48, Proposition 4.1.8 and Proposition 4.1.9]. This is obtained by assigning to an irreducible representations $\pi$ of $B$ an irreducible representations $\rho_{\pi}$ of $A$ such that $\pi$ is equivalent to the restriction of $\rho_{\pi}|_{B}$ to its essential subspace [48, Proposition 4.1.8]. Then we have that for every $a\in A$ and irreducible representation $\pi$ of $B$,
$w_{x^{\ast}ax}\left(\mathrm{\mathrm{Ker}}\left(\pi\right)\right)=\mathrm{\mathrm{rank}}\left(\pi\left(x^{\ast}ax\right)\right)\leq\mathrm{\mathrm{rank}}\left(\rho_{\pi}\left(x^{\ast}ax\right)\right)\leq\mathrm{\mathrm{rank}}\left(\rho_{\pi}\left(x\right)\right)=w_{x}\left(\mathrm{\mathrm{Ker}}\left(\rho_{\pi}\right)\right)\text{.}$
Since $x^{\ast}Ax$ is dense in $B$, this shows that, for every $b\in B$,
$w_{b}\left(\mathrm{\mathrm{Ker}}\left(\pi\right)\right)\leq w_{x}\left(\mathrm{\mathrm{Ker}}\left(\rho_{\pi}\right)\right)\text{.}$
Thus, by Lemma 5.8,
$r_{P}\left(b\right)\leq r_{P}\left(x\right)\text{.}$
As this holds for every $b\in B$, we conclude
$r_{P}\left(B\right)\leq r_{P}\left(x\right)\text{.}$
Consider now the inverse function
$X\setminus\mathrm{hull}\left(B\right)\rightarrow Y\text{, }\mathrm{\mathrm{Ker}}\left(\rho\right)\mapsto\mathrm{\mathrm{Ker}}\left(\rho_{B}\right)$
as in [48, Proposition 4.1.8 and Proposition 4.1.9]. This obtained by assigning to an irreducible representation $\rho$ of $A$ that does not vanish on $B$, the restriction of $\rho|_{B}$ on its essential subspace; see also [7, Proposition II.6.1.9]. Then we have that for $\mathrm{\mathrm{Ker}}\left(\rho\right)\in X\setminus\mathrm{hull}\left(B\right)$,
$w_{x}\left(\rho\right)=\mathrm{\mathrm{rank}}\left(\rho\left(x\right)\right)=\mathrm{\mathrm{rank}}\left(\rho_{B}\left(x\right)\right)=w_{x}\left(\rho_{B}\right)\text{.}$
Considering that, when $\mathrm{\mathrm{Ker}}\left(\rho\right)\in\mathrm{hull}\left(B\right)$, $w_{x}\left(\rho\right)=0$, we obtain from these remarks and Lemma 5.8 that
$r_{P}\left(x\right)\leq r_{P}\left(A\right)\text{,}$
concluding the proof. ∎
9. Catching them all
In this section we prove that any separable liminary C*-algebra is type $I_{\alpha}$ for some countable ordinal $\alpha$. Likewise, any separable C*-algebra with no infinite-dimensional representation is $\alpha$-subhomogeneous for some $\alpha<\omega_{1}$.
9.1. Characterization
We being with the characterization of $\alpha$-subhomogeneous C*-algebras.
Theorem 9.1.
Let $A$ be a separable C*-algebra. The following assertions are equivalent:
  1. (1)
    AA is preliminary;
  2. (2)
    AA is $\alpha$-subhomogeneous for some countable ordinal $\alpha$.
Proof.
[…]
$r_{P}\left(a\right)\leq r_{P}\left(A\right)\leq r_{P}\left(A^{\dagger}\right)=r_{P}\left(1_{A^{\dagger}}\right)<\omega_{1}$
by Proposition 7.3.
[…]
9.2. Liminary C*-algebras
Recall that a separable C*-algebra $A$ is liminary or CCR if for every irreducible representation $\pi$ of $A$, the range of $\pi$ is equal to the algebra of compact operators.
Theorem 9.2.
Let $A$ be a separable C*-algebra. The following assertions are equivalent:
  1. (1)
    AA is liminary;
  2. (2)
    AA is type $I_{\alpha}$ for some countable ordinal $\alpha$.
Proof.
Denote by $X$ the primitive spectrum of $A$.
[…]
$X_{\alpha}:=\left\{\mathrm{\mathrm{Ker}}\left(\pi\right)\in X:r_{F}\left(\pi\right)\leq\alpha\right\}$
is open in the Jacobson topology of $X$. The union of $X_{\alpha}$ for $\alpha<\omega_{1}$ is equal to $X$. By [34, Theorem 6.9], it follows that there exists $\alpha<\omega_{1}$ such that $X=X_{\alpha}$, and hence $r_{F}\left(A\right)\leq\alpha$.
[…]
10. Arbitrarily high rank
In this section we present, for any countable ordinal, examples of separable C*-algebra of Fell rank and Pedersen rank $\alpha$. For each limit ordinal $\lambda$ we fix an increasing sequence $\left(\lambda_{n}\right)$ of successor ordinals converging to $\lambda$. We also set $\alpha_{n}:=\alpha-1$ if $\alpha$ is a successor.
10.1. Bi-unitization
We say that a bipointed C*-algebra is a unital C*-algebra $A$ endowed with two distinguished characters $\infty$, $\infty^{\prime}$ and a projection $p$ such that $\infty\left(p\right)=\infty^{\prime}(1-p)=1$ and $A\cong B^{\dagger\dagger}$ where $B=\mathrm{\mathrm{Ker}}\left(\omega|_{pAp}\right)$, via an isomorphism that maps the distinguished characters and projection in $B^{\dagger\dagger}$ to $\infty$, $\infty^{\prime}$, and $p$ respectively.
In a similar fashion as in the case of the unitization, one can define a noncommutative analogue of the two-point compactification $X^{\dagger\dagger}$ of a countably compactly based $T_{0}$ space $X$ as considered in [37]. If $A$ is a C*-algebra, then we let $A^{\dagger\dagger}$ be the algebra $M_{2}\left(A\right)\oplus D_{2}\left(\mathbb{C}\right)$ where $D_{2}\left(\mathbb{C}\right)\subseteq M_{2}\left(\mathbb{C}\right)$ is the algebra of diagonal matrices, with coordinate-wise addition and scalar multiplication. The multiplication is induced by the $D_{2}\left(\mathbb{C}\right)$-bimodule structure on $M_{2}\left(A\right)$. Likewise, the C*-norm is the operator norm induced by letting $A^{\dagger\dagger}$ act as multiplication operators on $M_{2}\left(A\right)$. Equivalently, fixing a faithful nondegenerate representation $\subseteq B\left(H\right), AA^{\dagger\dagger}$ is the C*-subalgebra of $B\left(H\right)\otimes M_{2}$ generated by $A\otimes 1$ and $1\otimes D_{2}\left(\mathbb{C}\right)$. We let $\infty_{A}$ and $\infty_{A}^{\prime}$ be the distinguished characters of $A^{\dagger\dagger}$ corresponding to the points at infinity of $X^{\dagger\dagger}$, which we call characters at infinity. The projection $p_{A}\in A^{\dagger\dagger}$ correspond to a matrix unit in $D_{2}\left(\mathbb{C}\right)$. This turns $A^{\dagger\dagger}$ into a bipointed C*-algebra.
[…]
10.2. The Lazar jump and Lazar C*-algebras
[…]
We can use the jump construction to recursively define a unital preliminary C*-algebra $L_{\alpha}$ for every countable successor ordinal $\alpha$, called the Lazar algebra of rank $\alpha$. We being with $L_{0}:=\mathbb{C}$. For $\alpha>0$ we then set
$L_{\alpha}:=\mathrm{L}(A_{\left(\alpha-1\right)_{n}}\otimes M_{n}\left(\mathbb{C}\right))_{n\in\mathbb{N}}\text{.}$
Then it is proved as in [37] by induction that $L_{\alpha}$ is a separable liminary unital C*-algebra. It is also easily verified that $L_{\alpha}$ satisfies Fell’s condition of order $1$, i.e. it has Fell rank $1$. However, if $\infty_{\alpha}$ denotes one of the two characters at infinity of $L_{\alpha}$, then
$r_{P}\left(A\right)=r_{P}\left(1_{A}\right)=r_{d_{1_{A}}}\left(\infty_{\alpha}\right)=\omega\alpha+1\text{.}$
Thus, $L_{\alpha}\otimes M_{d}\left(\mathbb{C}\right)$ has Pedersen rank $\omega\alpha+d$.
10.3. The Taylor jump and Taylor C*-algebras
[…]
$\infty_{n}^{\prime}\left(x_{n}\right)=\infty_{n+1}\left(x_{n+1}\right)$
for every $n\in\omega$. Define then $\mathrm{J}\left(A_{n}\right)_{n\in\mathbb{N}}$ to be $B^{\dagger\dagger}$.
Iterating this construction, one define the Taylor algebra $K_{\alpha}$ for every countable successor ordinal $\alpha$. We set $K_{0}=\mathbb{C}$. For a successor ordinal $\alpha>0$ define
$K_{\alpha}:=\mathrm{J}(K_{\left(\alpha-1\right)_{n}})_{n\in\mathbb{N}}$
with distinguished characters $\infty_{\alpha}$ and $\infty_{\alpha}^{\prime}$. Again, we have that $K_{\alpha}$ is a separable unital liminary C*-algebra. However, in this case the character $\infty_{\alpha}$ of $K_{\alpha}$ has Fell rank $\omega\alpha+1$, and this maximizes the Fell ranks of irreducible representations of $K_{\alpha}$. Thus, $K_{\alpha}\otimes M_{d}\left(\mathbb{C}\right)$ has Fell rank $\omega\alpha+d$, which in this case is also equal to the Pedersen rank.